OpenStudy (anonymous):

Solve the triangle. Round lengths to the nearest tenth and angle measures to the nearest degree. a = 6, c = 11, B = 109°

OpenStudy (anonymous):

law of cosines for this one

OpenStudy (fibonaccichick666):

draw a picture to start

OpenStudy (anonymous):

make that \[b^2=a^2+c^2-2ac\cos(B)\]

OpenStudy (anonymous):

fill in the numbers and you get \(b^2\) right away with a calculator then use the law of sines to find the other angle

OpenStudy (anonymous):

this is what i get, very close to 200 for \(b\) http://www.wolframalpha.com/input/?i=6^2%2B11^2-2*6*11*cos%28109%29

OpenStudy (anonymous):

i mean for \(b^2\)

OpenStudy (anonymous):

These are my choices A. b = 14.1, A = 24°, C = 47° B. b = 19.9, A = 22°, C = 49° C. b = 17, A = 26°, C = 45° D. no triangle

OpenStudy (fibonaccichick666):

So, did you draw a picture?

OpenStudy (anonymous):

I'm not quite sure how to

OpenStudy (anonymous):



OpenStudy (fibonaccichick666):

Ok so draw a triangle that has a big 120 degree angle, one side that is 6, and another that is 11. Can you do that?

OpenStudy (anonymous):